OpenStudy (anonymous):

Solve for x. 2/x = 6/2 2/3 1 1/2 6

OpenStudy (anonymous):

@sleepyhead314

OpenStudy (sleepyhead314):

6/2 = ? simplify that first then multiply both sides by x

OpenStudy (anonymous):

3?

OpenStudy (sleepyhead314):

#thatawkwardmomentwhenyouhaveahighersmartscorethanmeforaskingtoomanyquestions yes so 2/x = 3 multiply both sides by x

OpenStudy (anonymous):

6?

OpenStudy (anonymous):

lol i suck at this:(

OpenStudy (sleepyhead314):

mmm like (2/x)*x = 3x becomes 2 = 3x then divide both sides by 3

OpenStudy (anonymous):

1?

OpenStudy (sleepyhead314):

2 = 3x /3 /3

OpenStudy (anonymous):

is the answer 1 and 1/2?

OpenStudy (sleepyhead314):

2 /3 = 2/3